REU 2008: SCHEDULE FOR WEEKS SIX AND SEVEN

Date  SESAME Babai  Lawler McReynolds/Schmidt
Mon, July 28 9–12, 1–4  4:15–5:30
Tue, July 29 9–12, 1–4 4:15–5:30  6:30–8:00
Wed, July 30 9–12, 1–4  4:15–5:30
Thu, July 31 9–12, 1–4 4:15–5:30  6:30–8:00
Fri, August 1 9–12, 1–4  4:15–5:30
Date  SESAME Babai  May  McReynolds/Schmidt
Mon, August 4 9–12, 1–4 4:15–5:30
Tue, August 5 9–12, 1–4 4:15–5:30 6:30–8:00
Wed, August 6 9–12, 1–4 4:15–5:30
Thu, August 7 9–12, 1–4 4:15–5:30 6:30–8:00
Fri, August 8 9–12, 1–4 4:15–5:30

Here is the schedule in summary form:

SIXTH WEEK:

SESAME: 9:00–12:00 and 1:00–4:00 M-F Babai (Discrete): 4:15–5:30 TTh, E133 Lawler (Probability): 4:15–5:30 MWF, E133 McReynolds and Schmidt (Geometry): TTh 6:30–8:00, E206

SEVENTH WEEK:

SESAME: 9:00–12:00 and 1:00–4:00 M-F Babai (Discrete): 4:15–5:30 MWF, R251 May (Finite Spaces): 4:15–5:30 TTh, E203 McReynolds and Schmidt (Geometry): TTh 6:30–8:00, E206